# and knot coloring invariants

### v2

@@ 8. Directions for future work
In light of our results, we propose questions for further research.
  1. (1)
[…]
@@
    In light of Proposition 7.2, do there exist Legendrian knots $\Lambda_{1}$ and $\Lambda_{2}$ and a medial GL-rack $M$ such that $\operatorname{Col}(\Lambda_{1},M)=\operatorname{Col}(\Lambda_{2},M)$ but $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda_{1}),M)\not\cong\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda_{2}),M)$? In the spirit of Corollary 7.4, do there exist such $\Lambda_{1}$ and $\Lambda_{2}$ whose classical invariants are identical? A positive answer would show that $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda),M)$ is a proper enhancement of $\operatorname{Col}(\Lambda,M)$ as an invariant of $\Lambda$.
@@
    Let $\mathbb{F}$ be a field, and let $M$ be a medial GL-rack. In 2023, Elhamdadi et al. [enhancements]*Theorems 4.2 and 5.1 properly enhanced medial quandle-valued invariants of smooth links using $\mathbb{F}$-algebra homomorphisms between quandle rings and colorings of smooth links by idempotents of quandle rings. Do similar proper enhancements of $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda),M)$ exist?
@@
    In light of Example 6.6, do tensor products of left-distributive racks, medial quandles (cf. [Hom]*Section 8.1), or Latin quandles have any interesting properties?
@@
    Theorem 6.9 shows that universal-algebraic tensor products make $\mathsf{Rack}$ and $\mathsf{GLR}$ into symmetric magmoidal categories with units, leading us to ask what their unital nuclei are; see [magmoidal]*Section 2.5. In particular, is $\otimes$ associative? If so, then $\otimes$ induces symmetric monoidal structures on $\mathsf{Rack}$ and $\mathsf{GLR}$.
@@
    Extend the results of Sections 36 and the algorithms in Appendix A to 4-Legendrian racks and 4-Legendrian biracks, which Kimura [leg-cocycle]*Section 4 introduced in 2024.
@@
    Transverse knots are knots that lie everywhere transverse to the standard contact structure on $\mathbb{R}^{3}$; see [etnyre]*Section 2.4. Can one define rack-theoretic invariants of transverse knots?
@@ A.1. Enumeration of small GL-racks
We obtained the numbers $g(n)$ from Algorithm A.1 and $g^{m}(n)$, $g_{q}(n)$, and $g^{m}_{q}(n)$ from Algorithm A.2. Meanwhile, the numbers $r(n)$, $r^{m}(n)$, $r_{q}(n)$, and $r^{m}_{q}(n)$ were originally computed by McCarron [rack-count] in 2010, Vojtěchovský and Yang [library] in 2019, Henderson et al. [quandle-count] in 2006, and Jedlička et al. [medial-quandles] in 2015, respectively. It appears that each of $g(n)$, $g^{m}(n)$, $g_{q}(n)$, and $g^{m}_{q}(n)$ in Table A.1 grows exponentially and at a much faster rate than its counterpart for classical racks.
@@ A.2. Classification of small GL-racks
We now distinguish the exhaustive search algorithms in GAP [GAP4] that we used to compute these isomorphism classes. We build upon the work of Vojtěchovský and Yang [library] in 2019, who classified racks up to order $11$ [library-site].
@@ A.2.1. Tabulation of GL-racks
On the author’s personal computer, running our implementation of Algorithm A.1 in GAP with $n=7$ took just under two minutes, while the computation for $n=8$ took seven days. By contrast, a similar program using Definition 3.9 took three hours for the $n=7$ computation. This shows that verifying Definition 3.1 is significantly easier than verifying the bi-Legendrian rack axioms.
Data: List isoClasses of isomorphism classes of GL-racks with underlying set $X=\{1,\dots,n\}$ returned by Algorithm A.1
Result: Lists $\mathcal{Q}_{n}$, $\mathcal{M}_{n}$, and $\mathcal{B}_{n}$ of isomorphism classes of GL-quandles, medial GL-racks, and medial GL-quandles, respectively, whose underlying set is $X$
begin
       $\mathcal{Q}_{n},\mathcal{M}_{n},\mathcal{B}_{n},\textnormal{{seen}}\leftarrow\emptyset$;
       foreach list $R=[s,\mathtt{u},\mathtt{d}]$ in isoClasses do
             if seen contains a list $[t,\textnormal{{isQuandle}},\textnormal{{isMedial}}]$ such that $t=s$ then
                   if isQuandle then Add($\mathcal{Q}_{n}$, $R$);
                   if isMedial then
                         Add($\mathcal{M}_{n}$, $R$);
                         if isQuandle then Add($\mathcal{B}_{n}$, $R$);
             else
                   if $\mathtt{d}=\mathtt{u}^{-1}$ then
                         isQuandle $\leftarrow$ true;
                         Add($\mathcal{Q}_{n}$, $R$);
                   else isQuandle $\leftarrow$ false;
                   isMedial $\leftarrow$ true;
                   foreach ordered triple $(x,y,z)$ in $X^{3}$ do
                         if $s_{s_{x}(z)}s_{y}\neq s_{s_{x}(y)}s_{z}$ then
                               isMedial $\leftarrow$ false;
                               break;
                  if isMedial then
                         Add($\mathcal{M}_{n}$, $R$);
                         if isQuandle then Add($\mathcal{B}_{n}$, $R$);
                  Add(seen,$[s,\textnormal{{isQuandle}},\textnormal{{isMedial}}]$);
Algorithm A.2 Classification of medial GL-racks, all GL-quandles, and medial GL-quandles of order $n$ up to isomorphism, given a classification of GL-racks of order $n$.
A.2.2. Tabulation of medial GL-racks and GL-quandles
Algorithm A.2 tests whether or not each GL-rack in the output of Algorithm A.1 is medial or a GL-quandle. To test for mediality, the algorithm simply verifies equation (1) for all elements $x,y,z\in X$. By Remark 3.10, to test whether a GL-rack $[s,\mathtt{u},\mathtt{d}]$ is a GL-quandle, it suffices to verify whether $\mathtt{d}=\mathtt{u}^{-1}$.
On the author’s personal computer, running our implementation of Algorithm A.2 in GAP with $n=7$ took just under an hour, while the $n=8$ computation took eight days.
A.3. Exhaustive searches for GL-rack coloring numbers
We now describe Algorithm A.3, which computes all colorings of the GL-rack of an oriented Legendrian link $\Lambda$ by each GL-rack in the list $\mathsf{isoClasses}$ computed by Algorithm A.1. Before running Algorithm A.3, the user must input a presentation of $\mathcal{G}(\Lambda)$ in terms of crossing relations between generators of $\mathcal{G}(\Lambda)$; see Section 7.3 for examples of such presentations. By the discussion in Section 7.3, it suffices for the algorithm to search for all valid solutions in $R$ to the inputted crossing relations.
Data: List isoClasses of isomorphism classes of GL-racks with underlying set $Y=\{1,\dots,n\}$ from Algorithm A.1 and a presentation of $\mathcal{G}(\Lambda)=[X,s^{\Lambda},\mathtt{u}^{\Lambda},\mathtt{d}^{\Lambda})]$
Result: List solutions whose elements are lists $[s,\mathtt{u},\mathtt{d},\mathbf{y}]$ such that the mapping $x_{i}\mapsto y_{i}$ defines a GL-rack homomorphism $\mathcal{G}(\Lambda)\to[Y,s,\mathtt{u},\mathtt{d}]$
begin
       $m\leftarrow|X_{\Lambda}|$;
[…]
       foreach GL-rack $[Y,s,\mathtt{u},\mathtt{d}]$ in isoClasses do
             foreach ordered $m$-tuple $\mathbf{y}\leftarrow(y_{1},\dots,y_{m})$ in $Y^{m}$ do
                   if all crossing relations are satisfied after replacing each $x_{i}\in X_{\Lambda}$, $s^{\Lambda}$, $\mathtt{u}^{\Lambda}$, and $\mathtt{d}^{\Lambda}$ with $y_{i}$, $s$, $\mathtt{u}$, and $\mathtt{d}$, respectively then Add(solutions, $[s,\mathtt{u},\mathtt{d},\mathbf{y}]$);
Algorithm A.3 Computation of colorings of an oriented Legendrian link $\Lambda$ by GL-racks of a given order $1\leq n\leq 11$ as computed by Algorithm A.1.
In particular, if $R=(Y,s,\mathtt{u},\mathtt{d})$ is a bi-Legendrian rack of order $n\leq 11$, then $\operatorname{Col}(\Lambda,R)$ is simply the number of lists in $\mathsf{solutions}$ produced by Algorithm A.3 whose first three list elements are $s$, $\mathtt{u}$, and $\mathtt{d}$. To distinguish two oriented Legendrian links $\Lambda_{1}$ and $\Lambda_{2}$, it suffices to run Algorithm A.3 twice, once inputting $\mathcal{G}(\Lambda_{1})$ and again inputting $\mathcal{G}(\Lambda_{2})$, and find a GL-rack $R$ in $\mathsf{isoClasses}$ such that $\operatorname{Col}(\Lambda_{1},R)\neq\operatorname{Col}(\Lambda_{2},R)$. For example, running Algorithm A.3 with $n=2,3$ is how we determined which GL-racks and homomorphisms to employ in Section 7.3. Running the algorithm with $n=5$ also gave us the following example.
Example A.1.
In this example, we use Algorithm A.3 to once again distinguish the Legendrian $6_{2}$ knots $\Lambda_{1}$ and $\Lambda_{2}$ on the left and right of Figure 5, respectively. This time, we use the $222$nd GL-rack $R$ of order $5$ listed in the data linked above, which is a Legendrian rack. Let $Y:=\{1,2,3,4,5\}$. In cycle notation, let $\sigma,\tau\in S_{5}$ be the $5$-cycles $\sigma:=(12345)$ and $\tau:=(13524)$. Let $R:=(Y,\sigma,\tau)_{\mathrm{perm}}$. By Corollary 3.13, $R$ is a Legendrian rack.
We input the relations of $\mathcal{G}(\Lambda_{1})$ in (9) and then those of $\mathcal{G}(\Lambda_{2})$ in (10) into our GAP implentation of Algorithm A.3. After running the program with $n=5$, the program outputs the text in Figure A.1 upon reaching $\mathsf{isoClasses}[222]=R$. The output states that $\operatorname{Col}(\Lambda_{1},R)=0\neq 5=\operatorname{Col}(\Lambda_{2},R)$, and the images of the generators $(x_{1},\dots,x_{7})$ of $\mathcal{G}(\Lambda_{2})$ under each element of $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda_{2}),R)$ are given by the orbit of $(1,2,3,5,1,4,5)\in Y^{7}$ under the action of the subgroup $\langle\sigma\rangle\leq S_{5}$ on $Y^{7}$.
In particular, Example A.1 yields an analogue of Corollary 7.4 for Legendrian racks.
Proposition A.2.
There exist Legendrian knots sharing the same classical invariants that are distinguished by the Legendrian rack coloring numbers originally defined in [ceniceros]*Proposition 1.
Finding all colorings of knot 1 by GL-rack 222 of 308…
Finding all colorings of knot 2 by GL-rack 222 of 308…
[ [(1,2,3,4,5), … (1,2,3,4,5)], (1,3,5,2,4), (1,3,5,2,4), 1, 2, 3, 5, 1, 4, 5 ]
[ [(1,2,3,4,5), … (1,2,3,4,5)], (1,3,5,2,4), (1,3,5,2,4), 2, 3, 4, 1, 2, 5, 1 ]
[ [(1,2,3,4,5), … (1,2,3,4,5)], (1,3,5,2,4), (1,3,5,2,4), 3, 4, 5, 2, 3, 1, 2 ]
[ [(1,2,3,4,5), … (1,2,3,4,5)], (1,3,5,2,4), (1,3,5,2,4), 4, 5, 1, 3, 4, 2, 3 ]
[ [(1,2,3,4,5), … (1,2,3,4,5)], (1,3,5,2,4), (1,3,5,2,4), 5, 1, 2, 4, 5, 3, 4 ]
Number of colorings of knot 1 by GL-rack 222 of 308: 0
Number of colorings of knot 2 by GL-rack 222 of 308: 5
Since their GL-rack coloring numbers are distinct, these knots are not Legendrian isotopic.
Figure A.1. Excerpt from the output of our GAP implementation of Algorithm A.3 with $n=5$. Here, knots $1$ and $2$ are the Legendrian knots in Figure 5, while GL-rack $222$ of $308$ is the Legendrian rack $R$ described in Example A.1.
@@ Appendix B Tabulation of GL-racks of orders 2, 3, and 4
For example, the $11$th entry in Table B.2 is the permutation GL-rack of order 3 with $\sigma=(123)$ and bi-Legendrian structure $[\mathtt{u},\mathtt{d}]=[(132),\operatorname{id}]$, which we used to prove Theorem 7.3.
@@
Table B.1. The four isomorphism classes of GL-racks of order 2.

### v3

@@ 7. Directions for future work
We conclude by proposing questions for further research. In the following, let $\mathcal{G}(\Lambda)$ denote the fundamental GL-rack of a Legendrian link $\Lambda$ (see [karmakar]).
@@
[…]
@@
    Let $\mathbb{F}$ be a field, and let $M$ be a medial GL-rack. In 2023, Elhamdadi et al. [enhancements]*Thms. 4.2 and 5.1 properly enhanced medial quandle-valued invariants of smooth links using $\mathbb{F}$-algebra homomorphisms between quandle rings and colorings of smooth links by idempotents of quandle rings. Do similar proper enhancements of $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda),M)$ exist?
@@
    In light of Example 6.6, do tensor products of left-distributive racks, medial quandles (cf. [Hom]*Sec. 8.1), or Latin quandles have any interesting properties?
@@
    Theorem 6.9 shows that universal-algebraic tensor products make $\mathsf{Rack}$ and $\mathsf{GLR}$ into symmetric magmoidal categories with units, leading us to ask what their unital nuclei are; see [magmoidal]*Sec. 2.5. In particular, is $\otimes$ associative? If so, then $\otimes$ induces symmetric monoidal structures on $\mathsf{Rack}$ and $\mathsf{GLR}$.
@@
    Extend the results of Sections 36 and the algorithms in Appendix A to 4-Legendrian racks and 4-Legendrian biracks, which Kimura [leg-cocycle]*Sec. 4 introduced in 2024.
@@
    Transverse knots are knots that lie everywhere transverse to the standard contact structure on $\mathbb{R}^{3}$; see [etnyre]*Sec. 2.4. Can one define rack-theoretic invariants of transverse knots?
@@ A.1. Enumeration of small GL-racks
We obtained the first four columns from Algorithm A.1, while the last four columns were originally computed by McCarron [rack-count] in 2010, Vojtěchovský and Yang [library] in 2019, Henderson et al. [quandle-count] in 2006, and Jedlička et al. [medial-quandles] in 2015, respectively. It appears that each of $g(n)$, $g^{m}(n)$, $g_{q}(n)$, and $g^{m}_{q}(n)$ in Table A.1 grows exponentially and at a much faster rate than its counterpart for classical racks.
@@ A.2. Classification of small GL-racks
We discuss the exhaustive search algorithms in GAP [GAP4] that we used to compute these isomorphism classes. We build upon the work of Vojtěchovský and Yang [library] in 2019, who classified racks up to order $11$ [library-site].
@@ A.2.1. Tabulation of GL-racks
On the author’s personal computer, running our implementation of Algorithm A.1 in GAP with $n=7$ took just under two minutes, while the computation for $n=8$ took about $100$ minutes. By contrast, a similar program using Definition 3.9 took three hours for the $n=7$ computation. This shows that verifying Definition 3.1 is significantly easier than verifying the bi-Legendrian rack axioms.
Our GAP implementation of Algorithm A.1 also tests whether the underlying rack of each GL-rack is medial, a quandle, or both. The former verifies equation (1), while the latter verifies the condition in Remark 3.10. We omit these obvious checks from the above pseudocode.
@@ Appendix B Tabulation of GL-racks of orders 2, 3, and 4
For example, the $11$th entry in Table B.2 is the permutation GL-rack of order 3 with $\sigma=(123)$ and bi-Legendrian structure $[\mathtt{u},\mathtt{d}]=[(132),\operatorname{id}]$.
Table B.1. The four isomorphism classes of GL-racks of order 2.
@@
Table B.2. The 13 isomorphism classes of GL-racks of order 3.